OpenStudy (anonymous):

What is the distance between the points (-3, -5) and (0, -1)?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

I got -25 is that correct?

hartnn (hartnn):

distance can never be negative how did you get -25 ?

OpenStudy (anonymous):

(-3-0)^2+(-5(-1))^2 (-3-0)^2+(-5+1) (-3)^2+(-4)^2 -9+-16=-25

hartnn (hartnn):

\((-3)^2= (-3)\times (-3)=9 \\ (-4)^2 =16\) got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so now whats the sum ?

hartnn (hartnn):

\(d =\sqrt{16+9}=...?\)

OpenStudy (anonymous):

25

hartnn (hartnn):

so, the distance d = \(\sqrt {25}=...?\)

OpenStudy (anonymous):

5

hartnn (hartnn):

thats correct! thats your distance :)

OpenStudy (anonymous):

thank you so much!

hartnn (hartnn):

welcome ^_^ and since you are new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)